Unit 21 Homework Assignment  Physics141-Summer 2009

Unit 21 homework after session 1

• Read chapters 14 and 25 in the text book Understanding Physics.

• Work textbook problems 14-12, 14-22, 14-31. Unit 21 homework after session 2

• Work textbook problems 25-6, 25-8, 25-25, 25-41.

All Unit 21 homework problems and the unit 21 activity guide are due on the first day of Unit 22.